OpenStudy (anonymous):

solve the equation. 16^3 = 4^x

OpenStudy (lgbasallote):

Hint: 16 = 4^2 another hint...a^x = a^y then x = y

OpenStudy (anonymous):

log(4) of both sides

OpenStudy (anonymous):

lgbasllote can help you visualize it, at the end you have 4^(2*3)=4^(x)